OpenStudy (anonymous):

Solve for x 83=27^x Consider that: log(a^x)=x (times) log(a)

OpenStudy (abb0t):

It looks like you already have your answer there. So just apply it and solve for x.

OpenStudy (mertsj):

\[\log(83)=xlog(27)\] \[x=\frac{\log(83)}{\log(27)}\]

OpenStudy (anonymous):

on the log(27) part, do you just ignore the x over the 27 ? or where did it goes ?

OpenStudy (mertsj):

That's where the rule comes in. if 8=2^3 then log8=3log2